Determine the oscillation frequency of the following circuit. (Choose the most-correct value) 1OK +6 Vo(t) 0.1AF TOK2 lOK2
Answer #1

Here Soo applid uoltage are our and by maximum output= 6+4= for minimum output =-4+6=24 This depending sipon different mode i